# INSIGHT: Universal Neural Simulator for Analog Circuits Harnessing Autoregressive Transformers

### v1

@@ Abstract.
Analog front-end design heavily relies on specialized human expertise and costly trial-and-error simulations, which motivated many prior works on analog design automation. However, efficient and effective exploration of the vast and complex design space remains constrained by the time-consuming nature of CPU-based SPICE simulations, making effective design automation a challenging endeavor. In this paper, we introduce INSIGHT, a GPU-powered, technology-independent, effective universal neural simulator in the analog front-end design automation loop. INSIGHT accurately predicts the performance metrics of analog circuits across various technology nodes, significantly reducing inference time. Notably, its autoregressive capabilities enable INSIGHT to accurately predict simulation-costly critical transient specifications leveraging less expensive performance metric information. The low cost and high fidelity feature make INSIGHT a good substitute for standard simulators in analog front-end optimization frameworks. INSIGHT is compatible with any optimization framework, facilitating enhanced design space exploration for sample efficiency through sophisticated offline learning and adaptation techniques. Our experiments demonstrate that INSIGHT-M, a model-based batch reinforcement learning framework that leverages INSIGHT for analog sizing, achieves at least 50×50$\times$50 × improvement in sample efficiency across circuits. To the best of our knowledge, this marks the first use of autoregressive transformers in analog front-end design.
@@ 1. Introduction
Analog circuits vary significantly across and within types, each characterized by distinct performance metrics, making analog front-end design challenging. The heavy reliance on human designer expertise and numerous rounds of costly trial-and-error simulations make the process labor-intensive. The consequent prolonged design cycles and increased time to market have motivated extensive prior research towards its automation.
@@
Various knowledge-based and optimization-based approaches have been proposed to automate analog front-end design, addressing circuit topology selection and device sizing. Knowledge-based methods, such as  1, 2, 3, 4, 5, are time-consuming and lack scalability. Optimization-based approaches, on the other hand, can be split into two categories: equation-based and simulation-based. Equation-based methods like  6, 7, 8, 9, 10, 11, 12 are highly inaccurate on modern technologies and lack scalability, while the simulation-based ones rely on expensive simulations. Evolutionary algorithms like Particle swarm optimization (PSO)  13, advanced differential evolution  14, and genetic algorithms  15, 16 require extensive SPICE simulations (sample inefficient). Bayesian optimization  17, 18 suffers from scalability issues arising from computational demands associated with model updates. Deep Reinforcement Learning (DRL) methods like  19, 20, 21, 22, 23, 24 demonstrate robustness but involve numerous simulator interactions. Modern algorithms  25, 26, 27 use black-box optimization and Model-Based Reinforcement Learning (RL) for device sizing. While they demonstrate improved sample efficiency, they still require many costly simulations. Given the vast and complex design space and the intricate trade-offs among performance metrics, repeated compute-intensive simulations within the optimization loops become inherent for effective and efficient design space exploration and exploitation to ensure optimal circuit performance. As a result, data-driven approaches have emerged as a promising solution, representing a recent trend in the field 28, 29.
Initially developed for language tasks 30, 31, transformer architectures 32 have been successfully extended to a diverse array of fields that involve recognizing complex patterns in the data, including vision 33, graph theory 34, and reinforcement learning 35, 36. This expansion was driven by the insight that many problems can be modeled as autoregressive sequence generation tasks. Autoregressive approaches significantly enhance the accuracy of output predictions, ensuring coherent, contextually aware outputs, compared to non-autoregressive approaches, with a minor trade-off in inference speed  37. Transformers have outperformed traditional models like convolutional 38, recurrent 39, and graph neural networks 40, demonstrating superior capabilities in handling complex data. Moreover, their ability to develop foundation models that learn generalized features from data, which can then be fine-tuned for specific tasks with exceptional performance, exemplified by GPT-3 30, has set new trends in AI. However, despite their ability to handle varying input and output sequence lengths, their application for building performance prediction models for analog circuits with diverse circuit parameters and specifications lengths remains unexplored.
@@
In this paper, we introduce INSIGHT, a novel GPU-powered technology-independent universal neural circuit simulator that accurately predicts performance on previously unseen parameters across various circuits and different technologies. Leveraging transformers’ inherent strengths, such as high adaptability and computational efficiency, INSIGHT consistently demonstrates high R2 scores and minimal test losses. Its autoregressive capabilities facilitate accurate prediction of simulation-costly critical transient specifications leveraging performance data from less expensive simulations. These features position INSIGHT as a cost-effective, high-fidelity alternative to standard circuit simulators in analog front-end optimization frameworks. INSIGHT is highly scalable and can seamlessly integrate into existing optimization workflows, facilitating computationally economical and constraint-free design space exploration. This integration significantly enhances traditional EDA optimization schemes, which often face bottlenecks due to computational constraints arising from the compute-intensive nature of simulations required at each design cycle iteration. Additionally, transformers’ robust transfer learning capabilities facilitate INSIGHT’s smooth adaptation to evolving technology landscapes while minimizing data requirements. We integrate INSIGHT within a Model-based Batch RL framework (INSIGHT-M) for circuit sizing and observe very high sample efficiencies for different circuits, underscoring the transformative capabilities of transformers in Analog Front-end Design Automation. To our knowledge, this marks the first application of transformers in Analog Front-end Design Automation.
We have validated our concept on different circuits covering different circuit types including operational amplifiers, trans-impedance amplifiers, comparators, and level shifters across different technology nodes.
@@
  • We introduce a novel approach to performance prediction for analog circuits by conceptualizing it as a sequence modeling problem. Inspired by recent advancements in natural language processing, we adapt transformer architectures, which are pivotal to modern language models, to decode intricate behaviors in circuits, thereby enhancing predictive accuracy and efficiency.
  • We propose INSIGHT, an innovative GPU-accelerated technology-independent universal neural circuit simulator. INSIGHT is designed to predict performance across diverse circuits and parameters, offering a versatile tool that significantly improves prediction speed and accuracy.
  • We have rigorously validated our approach across six different circuits across multiple technology nodes. INSIGHT consistently achieves high R2 scores and minimal loss scores across circuits and technology nodes. Its autoregressive capabilities enable accurate prediction of simulation-costly critical transient specifications leveraging less expensive performance metric information, thus demonstrating its capability as a high-fidelity substitute for standard simulators. By embedding INSIGHT within a Model-based Batch RL framework, INSIGHT-M, for circuit sizing optimization, we achieved exceptional sample efficiencies, with at least 50×50$\times$50 × improvement across circuits, thus underscoring INSIGHT’s potential to revolutionize diverse optimization frameworks.
  • To the best of our knowledge, this is the first application of transformer models in analog front-end design automation, paving the way for leveraging recent advancements in language models to enhance circuit design and optimization.
The rest of the paper is organized as follows. Section 2 describes the background and motivation; Section 3 explains the detailed implementation; Section 4 demonstrates the results; Section 5 concludes the paper.
@@ 2. Preliminaries
In this section, we delve into the fundamentals of analog circuit physics (Section  2.1) and explore transformer architectures (Section  2.2), discussing our rationale for choosing them. We also highlight the definition of a key design quality metric in Section  2.3, called Figure of Merit (FOM), which will guide our sizing flow.
2.1. Analog Circuit Physics
In analog circuits, performance metrics and parameters are highly interdependent. By leveraging these interdependencies, we can build effective prediction models. For instance, fundamental metrics such as Quiescent Current ($I_{Q}$) and DC Gain, when combined with parameter information, can significantly enhance predictions for other key specifications, including unity-gain bandwidth (UGBW) and phase margin (PM). Similarly, combining all these metrics alongside parameter information provides crucial insights into the circuit’s potential transient behavior. Thus, integrating partial performance information alongside parameter information can significantly enhance prediction accuracies for the other performance specifications, including critical transient specifications such as slew rate and settling time, which are typically costly and slow to simulate. This insight motivates the adoption of autoregressive models for precise performance prediction. These models leverage partial performance information along with parameter information to provide a richer context for predicting the remaining metrics, aiming to build a high-fidelity substitute for standard simulators. Furthermore, given the transferable nature of circuit physics across technology nodes, understanding performance interdependencies within one node provides valuable insights applicable to other nodes. Consequently, a model exhibiting high predictive accuracy for circuit performance on one technology node will likely achieve similar success on subsequent nodes and be easily transferable across advanced technology nodes. These insights, combined with the adaptability and transferability of transformer architectures, motivate the use of autoregressive transformers to build a universal neural simulator that would work efficiently across technologies.
@@ 2.2. Transformer Architectures
Transformers  32 constitute a powerful class of generative deep learning architectures that have revolutionized sequence modeling tasks in natural language processing (NLP) and beyond. These models, foundational for various state-of-the-art language models, stand out due to their self-attention mechanism, effectively capturing dependencies between tokens in a sequence, regardless of distance. The self-attention mechanism evaluates dependencies within a sequence by computing a weighted sum of value vectors. These weights highlight the most relevant token information for predicting subsequent tokens and are derived from the dot product of input vectors. Additionally, unlike traditional RNNs 41, transformers process data in parallel, significantly improving efficiency and scalability.
While the original Transformer architecture 32 consists of stacked layers of encoders and decoders, decoder-only models like GPT 42 have recently demonstrated notable success in sequence data generation. Decoder-only Transformers employ solely the decoder component, composed of multi-head self-attention and position-wise feed-forward networks. Positional encoding is also incorporated to retain the sequence order, crucial for maintaining context and coherence. This architectural choice is optimized for autoregressive tasks, where the model iteratively predicts the subsequent token by leveraging all previously generated tokens along with the input sequence as context.
From a probabilistic perspective, in the context of autoregressive generation, transformers model the joint probability of a sequence of tokens ($x_{1}$,$x_{2}$,…,$x_{N}$) by decomposing it into a product of conditional probabilities:
(1) $P(x_{1},x_{2},\dots,x_{N})=\prod_{i=1}^{N}P(x_{i}|x_{1},x_{2},\dots,x_{i-1})$
Each term $P(x_{i}|x_{1},x_{2},\dots,x_{i-1})$, which quantifies the likelihood of observing the token $x_{i}$, given the sequence of all preceding tokens ($x_{1},x_{2},\dots,x_{i-1}$), is estimated using the self-attention mechanism. The mechanism’s ability to capture long-range dependencies and contextual relationships within data, along with its adaptability to variable sequence lengths, makes Transformers ideal for handling complex sequences with widely varying input sizes. Drawing inspiration from this probabilistic formulation, we aim to enhance analog circuit performance prediction by strategically integrating partial performance data with parameter information, providing richer context as detailed in the following section.
@@ 2.3. Figure of Merit
Analog circuit sizing involves determining the optimal device sizes to meet specified performance criteria. Formally, given an objective performance metric $f_{0}(\mathbf{x})$ and $m$ constraint metrics $\{f_{i}(\mathbf{x})\leq 0\mid i=1,\ldots,m\}$, we adopt the Figure of Merit (FoM) definition given in  25, 43, 26, to evaluate quality of a design. This is of the following form:
@@
where $w_{i}$ is the weighting factor, a $\max(\cdot)$ function is used for equating designs after constraints are met, and $\min(\cdot)$ is used to prevent a single constraint violation from dominating the FoM value.
3. Algorithms
In this section, we introduce our proposed frameworks. We begin by detailing the construction of our proposed universal neural simulator for analog circuits, INSIGHT. Figure 1 illustrates how INSIGHT conceptualizes the prediction of performance metrics as an autoregressive sequence generation task, with Section 3.1 providing the probabilistic rationale for this approach. Section 3.2 then offers a detailed discussion of our analog circuit sizing optimization algorithm leveraging INSIGHT. As illustrated in Figure 2, the analog circuit sizing optimization algorithm for a given topology consists of two primary phases. The initial phase, discussed in Section 3.2.1, consists of a one-time effort for simulation-based performance dataset generation and training of INSIGHT. The second phase, discussed in Section 3.3, involves the integration of the one-time trained INSIGHT into a Model-Based Batch RL framework, termed INSIGHT-M, to perform automated analog sizing optimization for previously unseen user-input target specifications.
3.1. INSIGHT
Formally, for a given analog circuit topology, given a sequence of design parameters ($x_{1}$ to $x_{N}$), such as width-to-length ratios (W/L), our goal is to predict the corresponding performance specifications ($y_{1}$ to $y_{M}$) accurately. This predictive task is mathematically framed as maximizing the conditional probability $P(y_{1},y_{2},\ldots,y_{M}\mid x_{1},x_{2},\ldots,x_{N})$, which can be expressed as follows:
@@ 3.1. INSIGHT
This objective is equivalent to minimizing the Negative Log Likelihood (NLL), expressed as:
@@
This formulation mirrors the contextual approach used in autoregressive sequence generation tasks, such as those involving decoder-based transformers designed for enhanced sequence analysis. In analog circuits, where performance metrics are inherently interdependent, this formulation attempts to leverage these intricate relationships to predict the performance metrics for previously unseen input parameters accurately. Figure 1 presents a conceptual diagram illustrating how INSIGHT implements autoregressive modeling to predict performance metrics. The self-attention mechanism within the Transformer model enables it to adaptively focus on different parts of the input parameter sequence ($x_{1}$ to $x_{N}$) combined with the previously determined performance outputs ($y_{1}$ to $y_{i-1}$), depending on what is most relevant for predicting each specific output metric $y_{i}$. This capability is crucial for uncovering insights into the physics of the circuit, as it allows the model to learn and utilize the dependencies of one performance metric on another along with parameter information. Table 1 presents INSIGHT’s architecture details.
Figure 1. Autoregressive circuit performance prediction in INSIGHT
3.2. Analog Circuit Sizing Optimization
@@
Figure 2 showcases our sizing optimization algorithm. It comprises of two main sections: INSIGHT Pre-Training Phase (Section  3.2.1) and INSIGHT-M (Section  3.3). The core of RL in this paper is PPO 44. We basically use the hyper-parameters therein. Note that other RL algorithms can be employed with INSIGHT.
@@
  Activation GeLU
@@
Activation ReLU

### v2

@@ Abstract.
Analog front-end design heavily relies on specialized human expertise and costly trial-and-error simulations, which motivated many prior works on analog design automation. However, efficient and effective exploration of the vast and complex design space remains constrained by the time-consuming nature of SPICE simulations, making effective design automation a challenging endeavor. In this paper, we introduce INSIGHT, a GPU-powered, technology-agnostic, effective universal neural simulator in the analog front-end design automation loop. INSIGHT accurately predicts the performance metrics of analog circuits across various technologies with just a few microseconds of inference time. Notably, its autoregressive capabilities enable INSIGHT to accurately predict simulation-costly critical transient specifications leveraging less expensive performance metric information. The low cost and high fidelity feature make INSIGHT a good substitute for standard simulators in analog front-end optimization frameworks. INSIGHT is compatible with any optimization framework, facilitating enhanced design space exploration for sample efficiency through sophisticated offline learning and adaptation techniques. Our experiments demonstrate that INSIGHT-M, a model-based batch reinforcement learning sizing framework with INSIGHT as the accurate surrogate, only requires $<$ 20 real-time simulations with 100-1000x lower simulation costs and significant speedup over existing sizing methods.
@@ 1. Introduction
Analog circuits vary significantly across and within types, each characterized by distinct performance metrics, making analog front-end design challenging. This complexity necessitates a labor-intensive process reliant on designer expertise and costly trial-and-error simulations. The consequent prolonged design cycles and increased time to market have motivated extensive research towards its automation.
@@
Various knowledge-based and optimization-based approaches, including equation-based and simulation-based methods, have been proposed for analog design automation. While knowledge-based methods, such as  1, 2, 3, 4, are time-consuming and lack scalability, equation-based methods, like  5, 6, 7, 8, are also highly inaccurate on modern technologies. Conversely, simulation-based methods, while effective, depend on costly real-time simulations. Among these, evolutionary algorithms like particle swarm optimization  9, advanced differential evolution  10, and genetic algorithms  11, 12 require extensive SPICE simulations (sample inefficient). Deep Reinforcement Learning (RL) methods like  13, 14, 15, 16, 17, 18, 19, 20 while robust, also involve frequent simulator interactions. Meanwhile, Bayesian Optimization  21, 22 suffers from scalability issues due to computational demands associated with model updates. Modern algorithms  23, 24, 25 use black-box optimization and model-based RL to enhance sample efficiency, yet they still require numerous costly simulations. Given the vast and complex design space and intricate trade-offs among performance metrics, repeated compute-intensive simulations within the optimization loops become inherent for effective and efficient design space exploration and exploitation to ensure optimal circuit performance.
This heavy reliance on expensive real-time simulations motivates us to find an efficient and effective surrogate for the standard SPICE simulator to expedite analog design optimization tasks. The complexity of parameter-performance relationships in analog circuits renders simple DNNs inadequate, necessitating bigger, more complex models and extensive, costly EDA data while still resulting in suboptimal accuracies 25. Given the high interdependence of performance metrics inherent to analog circuits, we propose to formulate the analog circuit performance prediction problem as an autoregressive generation task, i.e., predicting each remaining simulation-costly performance specification by leveraging provided design parameters and partial performance information (provided or previously-generated) in a sequential manner. The autoregressive problem setup enforces the model to automatically reason and capture the inherent interdependency in the challenging analog performance prediction task, easing the model learning complexity, therefore, providing much better prediction quality than vanilla DNN setup 25.
@@
In this paper, we introduce INSIGHT, a novel, technology-agnostic, universal neural circuit simulator that accurately predicts performance on previously unseen parameters across various circuits and technologies autoregressively, consistently demonstrating R2 scores as high as 0.99 and minimal test losses at microsecond inference times. Its autoregressive capabilities facilitate accurate prediction of simulation-costly critical transient specifications leveraging performance data from less expensive simulations. These features position INSIGHT as a cost-effective and high-fidelity alternative to standard circuit simulators in analog front-end optimization frameworks. INSIGHT is highly scalable and can integrate into existing EDA optimization workflows to enhance efficiency. To showcase the effectiveness of INSIGHT as a surrogate model, we integrate INSIGHT within a model-based batch RL framework (INSIGHT-M) for circuit sizing and observe very high sample efficiencies for different circuits, covering several circuit types, like operational amplifiers, transimpedance amplifiers, comparators, and level shifters, across various technologies.
@@
  • We introduce a novel problem formulation for analog circuit performance prediction by conceptualizing it as an autoregressive sequence modeling problem, which significantly eases the learning complexity of the neural model.
  • We build INSIGHT, a GPU-accelerated, technology-agnostic universal neural simulator based on a decoder-only Transformer architecture to solve the analog circuit performance prediction problem. INSIGHT enables accurate prediction of simulation-costly critical transient specifications using less expensive performance metrics and and demonstrates its potential as a low-cost, high-fidelity surrogate model in the analog design automation loop.
  • We have rigorously validated our INSIGHT across various circuit and performance metric types across multiple technologies. We also demonstrate INSIGHT-M, which leverages INSIGHT as the performance predictor for the device sizing problem. INSIGHT-M achieved exceptional sample efficiencies on various circuit types, requiring $<$ 20 real-time simulations compared to existing sizing methods, which require 100-1000x more simulations.
  • To the best of our knowledge, this marks the first successful application of a universal neural simulator framework leveraging autoregressive Transformers for analog design automation.
@@ 2. Preliminaries
In this section, we briefly discuss some insights into analog circuit basics/metrics (Section  2.1) and Transformer architectures (Section  2.2), discussing our rationale for choosing them. Additionally, we define a key design quality metric, the Figure of Merit (FoM), which will guide our sizing flow (Section  2.3).
2.1. Analog Circuit Basics and Metrics
In analog circuits, performance metrics and parameters are highly interdependent. By leveraging these interdependencies, we can build effective prediction models. For instance, fundamental metrics such as Quiescent Current ($I_{Q}$) and DC Gain, when combined with parameter information, can significantly enhance predictions for other key specifications, like unity-gain bandwidth (UGBW) and phase margin (PM). Similarly, integrating all static metrics provides crucial insights into the circuit’s potential transient behavior. Thus, leveraging less expensive performance metric information can significantly enhance prediction accuracies for critical, simulation-costly, transient specifications such as slew rate and settling time. This insight motivates the adoption of autoregressive models for precise performance prediction. Furthermore, given the transferable nature of circuit physics across technologies for a given topology, understanding performance interdependencies within one node provides valuable insights applicable to other nodes. These insights, combined with the adaptability and transferability of Transformer architectures, motivate the adoption of autoregressive Transformers to build a universal neural simulator that would work efficiently across technologies.
@@ 2.2. Transformer Architectures
Transformers constitute a powerful class of generative deep learning architectures that have revolutionized sequence modeling tasks in natural language processing (NLP) and beyond. These models stand out due to their self-attention mechanism, which effectively captures dependencies between tokens in a sequence regardless of distance. The self-attention mechanism evaluates dependencies within a sequence by computing a weighted sum of value vectors. These weights highlight the most relevant token information for predicting subsequent tokens and are derived from the dot product of input vectors. Unlike traditional RNNs 26, Transformers process data in parallel, enhancing efficiency and scalability. While the original Transformer architecture  27 consists of stacked layers of encoders and decoders, decoder-only models like GPT 28 have recently demonstrated notable success in sequence data generation. Decoder-only Transformers employ solely the decoder component, composed of multi-head self-attention and position-wise feed-forward networks.
@@ 2.3. Figure of Merit
Given an objective performance metric $f_{0}(\mathbf{x})$ and $m$ constraint metrics $\{f_{i}(\mathbf{x})\leq 0\mid i=1,\ldots,m\}$, we adopt the Figure of Merit (FoM) definition given in  23, 29, 24, to evaluate quality of a design. This is of the following form:
@@
where $w_{i}$ is the weighting factor, a $\max(\cdot)$ function is used for equating designs after constraints are met, and $\min(\cdot)$ is used to prevent a single constraint violation from dominating the FoM value. This formulation will guide our sizing workflow.
3. INSIGHT Framework and Algorithms
In this section, we introduce our proposed frameworks. We begin by detailing the construction of INSIGHT, our proposed universal neural simulator for analog circuits, along with the probabilistic rationale for this approach, in Section 3.1. Section  3.2 then details our analog circuit sizing optimization framework for a given topology (Figure 2).
3.1. INSIGHT for Modeling
In INSIGHT, we are driven by a critical inquiry: Can we develop an effective and efficient surrogate model that utilizes circuit parameters and inexpensive partial performance metrics information (if available) to predict remaining simulation-costly performance numbers accurately? This question is particularly pertinent as current analog front-end design automation heavily depends on time-intensive simulations like SPICE, incurring a strong demand to ease the reliance.
The problem can be formally represented as follows: For a given analog circuit topology, given a sequence of design parameters ($x_{1}$ to $x_{N}$), such as width-to-length ratios (W/L), our goal is to predict the corresponding performance specifications ($y_{1}$ to $y_{M}$) accurately. This predictive task is mathematically framed as maximizing the conditional probability $P(y_{1},y_{2},\ldots,y_{M}\mid x_{1},x_{2},\ldots,x_{N})$, which can be decomposed into a product of conditional probabilities as follows:
@@ 3.1. INSIGHT for Modeling
Each term $P(y_{i}|x_{1},x_{2},\dots,x_{N},)$, which quantifies the likelihood of observing the token $y_{i}$, given the design parameters and the predicted performance specifications. This objective is equivalent to minimizing the Negative Log Likelihood (NLL), expressed as:
@@
This formulation connects to the contextual approach in autoregressive sequence generation tasks involving decoder-based Transformers like GPT 28 and motivates us to use a decoder-style Transformer model for INSIGHT. Figure 1 presents a conceptual diagram illustrating how INSIGHT implements autoregressive modeling to predict performance metrics. We utilize the self-attention mechanism to enable the model to adaptively focus on different parts of the input parameter sequence ($x_{1}$ to $x_{N}$) combined with the previously determined performance outputs ($y_{1}$ to $y_{i-1}$), depending on what is most relevant for predicting each specific output metric $y_{i}$. Positional encoding is also incorporated to retain the sequence order, crucial for maintaining context and coherence. The autoregressive feature ensures the model learns to utilize dependencies of one performance metric on another, along with the parameter information, to predict the performance metrics accurately. Table 1 presents INSIGHT’s architecture details.
Figure 1. Autoregressive Performance Prediction in INSIGHT
@@
  Activation GeLU 30
  Learning rate 0.001
with cosine annealing scheduler  31
@@
Activation ReLU 32
  Learning rate 0.001
Although our analog performance prediction task is analogous to the sequence generation tasks, we would like to emphasize that we need to predict real values instead of generating a sequence of tokens, such as a word or a position in an image, from a set of finite elements, or generate scalar real values from sequential contexts, as shown in previous Transformer works. Generating a sequence of real values based on Transformer architecture is unclear  33, and thus, suitable foundation models for predicting circuit parameters do not exist. Therefore, it is essential to evaluate the feasibility of Transformer architecture in generating a sequence of circuit outputs to confirm the potential for constructing circuit foundation models.
Accordingly, building on the existing decoder-Transformer architecture 28, 34, we show that our proposed architecture accelerates inference and training to demonstrate the feasibility of a unified framework like a foundation model. This streamlined approach enhances processing efficiency, reduces data requirements, and makes the model more practical for various applications. By focusing on simplicity, we improve performance and usability while maintaining robust predictions.
3.2. INSIGHT for Analog Sizing
Figure 2 showcases our sizing optimization algorithm. It comprises of two main sections: INSIGHT Pre-Training Phase (Section  3.2.1) and INSIGHT for Optimization (Section  3.2.2).